# Text Indexing and Pattern Matching with Ephemeral Edits

### v1

@@ Abstract
We introduce text indexing with ephemeral substring edits, a new version of text indexing. Our goal is to design a data structure over a given text that supports subsequent pattern matching queries with ephemeral substring insertions, deletions, or substitutions in the text; we require insertions and substitutions to be of constant length. In particular, we preprocess a text $T=T[0\mathinner{.\,.}n)$ over an integer alphabet $\Sigma=[0,\sigma)$ with $\sigma=n^{\mathcal{O}(1)}$ in $\mathcal{O}(n)$ time. Then, we can preprocess any arbitrary pattern $P=P[0\mathinner{.\,.}m)$ given online in $\mathcal{O}(m\log\log m)$ time and $\mathcal{O}(m)$ space and allow any ephemeral sequence of edit operations in $T$. Before reverting the $i$th operation, we report all Occ occurrences of $P$ in $T^{i}$ in $\mathcal{O}(\log\log n+\textsf{Occ})$ time.
We also introduce pattern matching with ephemeral edits, a new version of pattern matching. Here, the goal is to design a data structure over two given strings that supports pattern matching with ephemeral edits in one of the two strings. In particular, we preprocess two strings $T$ and $P$, each of length at most $n$, over an integer alphabet $\Sigma=[0,\sigma)$ with $\sigma=n^{\mathcal{O}(1)}$ in $\mathcal{O}(n)$ time. Then, we allow any ephemeral sequence of edit operations in $T$. Before reverting the $i$th operation, we report all Occ occurrences of $P$ in $T^{i}$ in the optimal $\mathcal{O}(\textsf{Occ})$ time. Along our way to this result, we also give an optimal solution for pattern matching with ephemeral block deletions.
@@ 1 Introduction
Pattern matching and text indexing are the two flagship problems of combinatorial pattern matching. In pattern matching, we are given a pattern $P=P[0\mathinner{.\,.}m)$ of length $m$ and a text $T=T[0\mathinner{.\,.}n)$ of length $n$, and we are asked to find all occurrences of $P$ in $T$. This can be solved in $\mathcal{O}(n)$ time by the classical Knuth-Morris-Pratt algorithm [KJP77]. In text indexing, we are asked to preprocess a text $T$ of length $n$ into a data structure that supports subsequent pattern matching queries: given a pattern $P$, we report (or count) all occurrences of $P$ in $T$. Reporting all Occ occurrences of $P$ in $T$ can be done in the optimal $\mathcal{O}(m+\textsf{Occ})$ time using suffix trees, which can be constructed in $\mathcal{O}(n)$ time [FFM00].
@@
In a typical setting, one already has at hand a long reference string as the text $T$ (e.g., a reference genome, a baseline legal document or a configuration file), which will remain mostly static, except for a few local edits. These local edits are often independent (e.g., genomic variants coming from individuals or client-specific variations); or hypothetical (e.g., they are not intended to be permanent, but rather are applied temporarily to test a specific scenario in a scientific simulation or to analyze the impact of a potential alteration). More concretely, this setting is highly relevant in computational pangenomics [Con18, TM20, MWHBS24], where $T$ models an existing reference sequence, and the sequence of independent edit operations represents a collection of short local variants originating from different individuals. [^1] The Variant Call Format (https://en.wikipedia.org/wiki/Variant_Call_Format) stores this type of data. In fact, many international consortia (e.g., https://alpaca-itn.eu/ and https://www.pangenome.eu/) deal with such data. The fact that local variants originate from different individuals suggests that once a variant is processed, the corresponding edit should be reverted before processing the next one. Indeed, when querying a pattern against one individual’s genome (which is a variant of the reference), we do not want that variant to persist when moving on to query against another individual’s genome, which will have its own distinct short local variant from the reference.
@@
We introduce a new interesting version of text indexing that we call text indexing with ephemeral substring edits. The goal is to design a data structure over a given text that supports subsequent pattern matching queries with ephemeral substring insertions, deletions, or substitutions in the text; we require insertions and substitutions to be short (of constant length), but deletions can be arbitrarily long. First, we preprocess a text $T$ of length $n$. Then, we are asked to preprocess any arbitrary pattern $P$ of length $m$ given online, and subsequently allow any ephemeral sequence of edit operations on $T$. Before reverting the $i$th operation, we must report all occurrences of $P$ in $T^{i}$. We show that after an $\mathcal{O}(n)$-time preprocessing of $T$, we can preprocess any arbitrary pattern $P$ of length $m$ in $\mathcal{O}(m\log\log m)$ time and $\mathcal{O}(m)$ space to then support pattern matching queries after any edit in $\mathcal{O}(\log\log n+\textsf{Occ})$ worst-case time. Note that $P$ is not static; we can preprocess any subsequent pattern in $\mathcal{O}(m\log\log m)$ time and $\mathcal{O}(m)$ space (i.e., independently of $n$). Further note that our query time is essentially independent of $m$. More formally, we show the following result.
@@
For any pattern $P$ of length $m$ and any constant $\epsilon\geq 1$ given online, we can preprocess $P$ in $\mathcal{O}(m\log\log m)$ time and $\mathcal{O}(m)$ space to support any sequence of the following queries:
@@
    Insert string $S$ of length $|S|\leq\epsilon$ after position $p$ of $T$, or delete fragment $T[q\mathinner{.\,.}p]$, for any $0\leq q\leq p<n$, or substitute fragment $T[p\mathinner{.\,.}p+|S|-1]$ by string $S$ of length $|S|\leq\epsilon$, for any $0\leq p\leq n-|S|$, to obtain text $T^{\prime}$ in $\mathcal{O}(1)$ worst-case time.
@@
Our technique consists in a novel $\mathcal{O}(m\log\log m)$-time preprocessing of $P$ into a collection of trees of $\mathcal{O}(m)$ total size. This preprocessing allows us to deploy predecessor search [NR21] and optimal prefix-suffix queries [Pis25, Gaw13] to process each edit in near-optimal time. With Theorem 1, we can support efficient pattern matching queries while processing ephemeral sequences of edits in $T$.
We also introduce a new version of pattern matching that we call pattern matching with ephemeral block deletions. The goal is to design a data structure over two given strings $T$ and $P$, each of length at most $n$, that supports pattern matching with ephemeral block deletions in $T$. First, we preprocess $T$ and $P$ in $\mathcal{O}(n)$ time. Then, we allow any ephemeral sequence of block deletions in $T$. Before reverting the $i$th operation, we report all occurrences of $P$ in $T^{i}$ in the optimal $\mathcal{O}(\textsf{Occ})$ worst-case time.
@@
Given a text $T$ and a pattern $P$, each of length at most $n$, over an integer alphabet $\Sigma=[0,\sigma)$ with $\sigma=n^{\mathcal{O}(1)}$, we can preprocess $T$ and $P$ in $\mathcal{O}(n)$ time and space to support any sequence of the following queries:
@@
    Delete $T[q\mathinner{.\,.}p]$, for any $0\leq q\leq p<n$, to obtain text $T^{\prime}$ in $\mathcal{O}(1)$ worst-case time.
@@
Our data structure relies on suffix trees [FFM00] and optimal prefix-suffix queries.
Finally, we introduce another version of pattern matching that we call pattern matching with ephemeral edits. The goal is to design a data structure over two given strings $T$ and $P$, each of length at most $n$, that supports pattern matching with ephemeral edit operations in $T$. First, we preprocess $T$ and $P$ in $\mathcal{O}(n)$ time. Then, we allow any ephemeral sequence of edit operations in $T$. Before reverting the $i$th operation, we report all occurrences of $P$ in $T^{i}$ in the optimal $\mathcal{O}(\textsf{Occ})$ worst-case time.
@@
Given a text $T$ and a pattern $P$, each of length at most $n$, over an integer alphabet $\Sigma=[0,\sigma)$ with $\sigma=n^{\mathcal{O}(1)}$, we can preprocess $T$ and $P$ in $\mathcal{O}(n)$ time and space to support any sequence of the following queries:
@@
    Insert letter $c$ after position $p$ of $T$, or delete letter $T[p]$, or substitute $T[p]$ by letter $c$ to obtain text $T^{\prime}$ in $\mathcal{O}(1)$ worst-case time.
@@
Our data structure relies on a somewhat novel combination of suffix trees and string matching automata [CHL07] as well as on optimal prefix-suffix queries.
The algorithms underlying Theorems 1, 2 and 3 are also simple to implement: they utilize only textbook algorithms or data structures with mature implementations; e.g., predecessor search [DFH21], suffix trees [GBMP14], string matching automata [CHL07], and range data structures [GBMP14].
@@
To the best of our knowledge, the existing dynamic text indexing algorithms have seen limited to no adoption by practitioners. One reason for this is certainly their extreme intricacy, as exemplified in [LMN24]. Another related reason is that, while a fully dynamic model is theoretically appealing, it often proves overly powerful for the above problems. Let us review the main works on the topic before we provide more concrete arguments to answer the above question. The problem of dynamic text indexing was introduced by Gu, Farach, and Beigel in 1994 [GFB94]. The problem asks to design a data structure (a dynamic index) that supports updates (insertions or deletions) to a text $T$ and pattern matching queries. The solution of [GFB94] (see also [AB20] for an amortized analysis) achieves $\mathcal{O}(\log n)$-time updates to $T$ and $\mathcal{O}(m+\textsf{Occ}\log i+i\log m)$-time queries, where $i$ is the number of edit operations executed so far. Ferragina [Fer97] later showed how to support insertions and deletions of text blocks with the query time being roughly proportional to the number of updates made so far. Soon after that, Ferragina and Grossi [FG99] improved this to support updates in $\mathcal{O}(\sqrt{n})$ time and queries in the optimal $\mathcal{O}(m+\textsf{Occ})$ time. Sahinalp and Vishkin [SV96] showed the first polylogarithmic data structure that supports $\mathcal{O}(\log^{3}n)$-time updates and $\mathcal{O}(m+\textsf{Occ}+\log n)$-time queries. The update time was later improved by Alstrup, Brodal, and Rauhe [ABR00] to $\mathcal{O}(\log^{2}n\log\log n\log^{\star}n)$ at the expense of slightly slower queries with $\mathcal{O}(m+\textsf{Occ}+\log n\log\log n)$ time. Finally, this was improved by Gawrychowski et al. [GKK+15] to $\mathcal{O}(\log^{2}n)$-time updates and $\mathcal{O}(m+\textsf{Occ})$-time queries.
@@
Say that we want to perform $k$ independent or hypothetical edits on an existing text $T$ of length $n$ and report all occurrences of a pattern $P$ of length $m$ after each edit. If we use the state-of-the-art algorithm for fully dynamic text indexing [GKK+15] as a black box, the total time will be $\mathcal{O}((n+k)\log^{2}(n+k)+km+\textsf{Output})$. Using Theorem 1 instead, the total time will be $\mathcal{O}(n+m\log\log m+k\log\log n+\textsf{Output})$. Note that, for such a task, our improvement is substantial: even if we ignore the polylogarithmic terms, we replace the $km$ additive term by $k$.
Amir et al. [ALLS07] have also considered a restricted version of dynamic pattern matching, where the pattern $P$ is static (i.e., it is always the same) and only letter substitutions are supported in $T$. The preprocessing time of their data structure is $\mathcal{O}(n\log\log m+m\sqrt{\log m})$ and the space used is $\mathcal{O}(n+m\sqrt{\log m})$. After each text update, the algorithm deletes all previous occurrences of the pattern that no longer match, and reports all new occurrences of the pattern in the text in $\mathcal{O}(\log\log m)$ time.
@@
Say that we want to perform $k$ independent or hypothetical substitutions on an existing text $T$ of length $n$ and report all occurrences of a pattern $P$ of length $m$ after each edit. If we use the algorithm for a dynamic text and a static pattern [ALLS07], the total time and space will be $\mathcal{O}(n\log\log m+m\sqrt{\log m}+k\log\log m+\textsf{Output})$ and $\mathcal{O}(n+m\sqrt{\log m})$, respectively. Using Theorem 3 instead, the total time and space will be $\mathcal{O}(n+k+\textsf{Output})$ and $\mathcal{O}(n)$, respectively.
Although ephemeral edits have already been used for many other string problems, such as longest common substrings [ACI+17, ACPR20], palindrome substrings [FNI+18, FNI+19, FM21, FNI+21], Lyndon substrings [UNI+18], and substring covers [MMSH24], such a model has not been defined for text indexing or pattern matching. We were thus also motivated to close this (surprising) gap in theory.
@@
In Section 2 we provide the necessary definitions and notation, as well as a few important tools that we use throughout. In Section 3, we present our solution to text indexing with ephemeral substring edits. In Section 4, we present our solution to pattern matching with ephemeral block deletions. In Section 5, we present our solution to pattern matching with ephemeral edits. We conclude this paper in Section 6 with some future proposals.

### v2

@@ Abstract
We introduce text indexing with ephemeral substring edits, a new version of text indexing. The goal is to design a data structure over a given text that supports subsequent pattern matching queries with ephemeral substring insertions, deletions, or substitutions in the text; we require insertions and substitutions to be of constant length. In particular, we preprocess a text $T=T[0\mathinner{.\,.}n)$ over an integer alphabet $\Sigma=[0,\sigma)$ with $\sigma=n^{\mathcal{O}(1)}$ in $\mathcal{O}(n)$ time. Then, we preprocess any arbitrary pattern $P=P[0\mathinner{.\,.}m)$ given online in $\mathcal{O}(m\log\log m)$ time and $\mathcal{O}(m)$ space and support any ephemeral sequence of edit operations in $T$. Before reverting the $i$th operation, we report all Occ occurrences of $P$ in $T^{i}$ in $\mathcal{O}(\log\log n+\textsf{Occ})$ time.
We also introduce pattern matching with ephemeral substring edits, a new version of pattern matching. The goal is to design a data structure over two given strings that supports pattern matching with ephemeral substring insertions, deletions, or substitutions in one of the two strings; we require insertions and substitutions to be of constant length. In particular, we preprocess two strings $T$ and $P$, each of length at most $n$, over an integer alphabet $\Sigma=[0,\sigma)$ with $\sigma=n^{\mathcal{O}(1)}$ in $\mathcal{O}(n)$ time. Then, we support any ephemeral sequence of edit operations in $T$. Before reverting the $i$th operation, we report all Occ occurrences of $P$ in $T^{i}$ in the optimal $\mathcal{O}(\textsf{Occ})$ time. Along our way to this result, we also give an optimal solution for pattern matching with ephemeral block deletions.
@@ 1 Introduction
Pattern matching and text indexing are the two flagship problems of combinatorial pattern matching. In pattern matching, we are given a pattern $P=P[0\mathinner{.\,.}m)$ of length $m$ and a text $T=T[0\mathinner{.\,.}n)$ of length $n$, and we are asked to find all occurrences of $P$ in $T$. This can be solved in the optimal $\mathcal{O}(n)$ time using the classical Knuth-Morris-Pratt algorithm [KJP77]. In text indexing, we are asked to preprocess a text $T=T[0\mathinner{.\,.}n)$ into a data structure that supports subsequent pattern matching queries: given a pattern $P=P[0\mathinner{.\,.}m)$, we report (or count) all occurrences of $P$ in $T$. Reporting all Occ occurrences of $P$ in $T$ can be done in the optimal $\mathcal{O}(m+\textsf{Occ})$ time using suffix trees, which can be constructed in $\mathcal{O}(n)$ time [FFM00].
@@
In a typical setting, one already has at hand a long reference string as the text $T$ (e.g., a reference genome, a baseline legal document or a configuration file), which will remain mostly static, except for a few local edits. These local edits are often independent (e.g., genomic variants coming from individuals or client-specific variations); or hypothetical (e.g., they are not intended to be permanent, but rather are applied temporarily to test a specific scenario in a scientific simulation or to analyze the impact of a potential alteration). More concretely, this setting is highly relevant in computational pangenomics [THE18, TM20, MWB+24], where $T$ models an existing reference sequence, and the sequence of independent edit operations represents a collection of short local variants originating from different individuals. [^1] The Variant Call Format (https://en.wikipedia.org/wiki/Variant_Call_Format) stores this type of data. In fact, many international consortia (e.g., https://alpaca-itn.eu/ and https://www.pangenome.eu/) deal with such data. The fact that local variants originate from different individuals suggests that once a variant is processed, the corresponding edit should be reverted before processing the next one. Indeed, when querying a pattern against one individual’s genome (which is a variant of the reference), we do not want that variant to persist when moving on to query against another individual’s genome, which will have its own distinct short local variant from the reference.
@@
We introduce a new interesting version of text indexing that we call text indexing with ephemeral substring edits. The goal is to design a data structure over a given text that supports subsequent pattern matching queries with ephemeral substring insertions, deletions, or substitutions in the text; we require insertions and substitutions to be short (of constant length), but deletions can be arbitrarily long. First, we preprocess a text $T$. Then, we are asked to preprocess any arbitrary pattern $P$ given online, and subsequently allow any ephemeral sequence of edit operations in $T$. Before reverting the $i$th operation, we must report all occurrences of $P$ in $T^{i}$. We show that after an $\mathcal{O}(n)$-time preprocessing of text $T$ of length $n$, we can preprocess any arbitrary pattern $P$ of length $m$ in $\mathcal{O}(m\log\log m)$ time and $\mathcal{O}(m)$ space to then support pattern matching queries after any edit in $\mathcal{O}(\log\log n+\textsf{Occ})$ worst-case time. Note that $P$ is not static; we can preprocess any subsequent pattern in $\mathcal{O}(m\log\log m)$ time and $\mathcal{O}(m)$ space (i.e., independently of $n$). Further note that our query time is essentially independent of $m$. More formally, we show the following result.
@@
For any pattern $P$ of length $m$ and any constant $\epsilon\geq 1$ given online, we can preprocess $P$ in $\mathcal{O}(m\log\log m)$ time and $\mathcal{O}(m)$ space to support the following sequence of queries:
@@
    Insert string $S$ of length $|S|\leq\epsilon$ after position $p$ in $T$, for any $1\leq p<n$, or delete fragment $T[q\mathinner{.\,.}p]$, for any $0\leq q\leq p<n$, or substitute fragment $T[p\mathinner{.\,.}p+|S|-1]$ by string $S$ of length $|S|\leq\epsilon$, for any $0\leq p\leq n-|S|$, to obtain text $T^{\prime}$ in $\mathcal{O}(1)$ worst-case time.
@@
Our technique underlying Theorem 1 consists in a novel $\mathcal{O}(m\log\log m)$-time preprocessing of $P$ into a collection of trees of $\mathcal{O}(m)$ total size. This preprocessing allows us to deploy predecessor search [NR21] and optimal prefix-suffix queries [PIS25, GAW13] to process each edit in near-optimal time. With Theorem 1, we can support efficient pattern matching queries while processing ephemeral sequences of edits in $T$.
We also introduce a new version of pattern matching that we call pattern matching with ephemeral block deletions. The goal is to design a data structure over two given strings $T$ and $P$, each of length at most $n$, that supports pattern matching with ephemeral block deletions in $T$. First, we preprocess $T$ and $P$ in $\mathcal{O}(n)$ time. Then, we support any ephemeral sequence of block deletions in $T$. Before reverting the $i$th operation, we report all occurrences of $P$ in $T^{i}$ in the optimal $\mathcal{O}(\textsf{Occ})$ worst-case time. More formally, we show the following result.
@@
Given a text $T$ and a pattern $P$, each of length at most $n$, over an integer alphabet $\Sigma=[0,\sigma)$ with $\sigma=n^{\mathcal{O}(1)}$, we can preprocess $T$ and $P$ in $\mathcal{O}(n)$ time and space to support the following sequence of queries:
@@
    Delete $T[q\mathinner{.\,.}p]$, for any $0\leq q\leq p<|T|$, to obtain text $T^{\prime}$ in $\mathcal{O}(1)$ worst-case time.
@@
Our data structure underlying Theorem 2 relies on suffix trees [FFM00] and optimal prefix-suffix queries.
Finally, we introduce another version of pattern matching that we call pattern matching with ephemeral substring edits. The goal is to design a data structure over two given strings $T$ and $P$, each of length at most $n$, that supports pattern matching with ephemeral substring insertions, deletions, or substitutions in $T$; we require insertions and substitutions to be short (of constant length), but deletions can be arbitrarily long, as indeed we rely on Theorem 2. First, we preprocess $T$ and $P$ in $\mathcal{O}(n)$ time. Then, we support any ephemeral sequence of edit operations in $T$. Before reverting the $i$th operation, we report all occurrences of $P$ in $T^{i}$ in the optimal $\mathcal{O}(\textsf{Occ})$ worst-case time. More formally, we show the following result.
@@
Given a text $T$ and a pattern $P$, each of length at most $n$, over an integer alphabet $\Sigma=[0,\sigma)$ with $\sigma=n^{\mathcal{O}(1)}$, we can preprocess $T$ and $P$ in $\mathcal{O}(n)$ time and space to support the following sequence of queries:
@@
    Insert string $S$ of length $|S|=\mathcal{O}(1)$ after position $p$ in $T$, for any $1\leq p<|T|$, or delete fragment $T[q\mathinner{.\,.}p]$, for any $0\leq q\leq p<|T|$, or substitute fragment $T[p\mathinner{.\,.}p+|S|-1]$ by string $S$ of length $|S|=\mathcal{O}(1)$, for any $0\leq p\leq|T|-|S|$, to obtain text $T^{\prime}$ in $\mathcal{O}(1)$ worst-case time.
@@
Our data structure underlying Theorem 3 relies on a somewhat novel combination of suffix trees and string matching automata [CHL07] as well as on optimal prefix-suffix queries.
The algorithms underlying Theorems 1, 2 and 3 are also simple to implement: they utilize only textbook algorithms or data structures with mature implementations; e.g., predecessor search [DFH21], suffix trees [GBM+14], string matching automata [CHL07], and range data structures [GBM+14].
@@
To the best of our knowledge, the existing dynamic text indexing algorithms have seen limited to no adoption by practitioners. One reason for this is certainly their extreme intricacy, as exemplified in [LMN24]. Another related reason is that, while a fully dynamic model is theoretically appealing, it often proves overly powerful for the above problems. Let us review the main works on the topic before we provide more concrete arguments to answer the above question. The problem of dynamic text indexing was introduced by Gu, Farach, and Beigel in 1994 [GFB94]. The problem asks to design a data structure (a dynamic index) that supports updates (insertions or deletions) to a text $T$ and subsequent pattern matching queries. The solution of [GFB94] (see also [AB20] for an amortized analysis) achieves $\mathcal{O}(\log n)$-time updates to $T$ and $\mathcal{O}(m+\textsf{Occ}\log i+i\log m)$-time queries, where $i$ is the number of edit operations executed so far and $m$ is the length of the pattern. Ferragina [FER97] later showed how to support insertions and deletions of text blocks with the query time being roughly proportional to the number of updates made so far. Soon after that, Ferragina and Grossi [FG99] improved this to support updates in $\mathcal{O}(\sqrt{n})$ time and queries in the optimal $\mathcal{O}(m+\textsf{Occ})$ time. Sahinalp and Vishkin [SV96] showed the first polylogarithmic data structure that supports $\mathcal{O}(\log^{3}n)$-time updates and $\mathcal{O}(m+\textsf{Occ}+\log n)$-time queries. The update time was later improved by Alstrup, Brodal, and Rauhe [ABR00] to $\mathcal{O}(\log^{2}n\log\log n\log^{\star}n)$ at the expense of slightly slower queries with $\mathcal{O}(m+\textsf{Occ}+\log n\log\log n)$ time. Finally, this was improved by Gawrychowski et al. [GKK+15, GKK+18] to $\mathcal{O}(\log^{2}n)$-time updates and $\mathcal{O}(m+\textsf{Occ})$-time queries. More recently, Kempa and Kociumaka [KK22] proposed a dynamic version of suffix arrays [MM93]—which is strictly harder than dynamic text indexing—supporting $\mathcal{O}(\log^{4}n)$-time updates.
@@
Say that we want to perform $k$ independent or hypothetical edits on an existing text $T$ of length $n$ and report all occurrences of a pattern $P$ of length $m$ after each edit. If we use the state-of-the-art algorithm for fully dynamic text indexing [GKK+15] as a black box, the total time will be $\mathcal{O}((n+k)\log^{2}(n+k)+km+\textsf{Output})$. Using Theorem 1 instead, the total time will be $\mathcal{O}(n+m\log\log m+k\log\log n+\textsf{Output})$. Note that our improvement is substantial: even if we ignore the polylogarithmic factors, we replace the $km$ term by $k$.
Amir et al. [ALL+07] have also considered a restricted version of dynamic pattern matching, where the pattern $P$ is static (i.e., it is always the same) and only letter substitutions are supported in $T$. The preprocessing time of their data structure is $\mathcal{O}(n\log\log m+m\sqrt{\log m})$ and the space used is $\mathcal{O}(n+m\sqrt{\log m})$. After each text update, the algorithm deletes all previous occurrences of the pattern that no longer match, and reports all new occurrences of the pattern in the text in $\mathcal{O}(\log\log m)$ time.
@@
Say that we want to perform $k$ independent or hypothetical substitutions on an existing text $T$ of length $n$ and report all occurrences of a pattern $P$ of length $m$ after each edit. If we use the algorithm for a dynamic text and a static pattern [ALL+07], the total time and space will be $\mathcal{O}(n\log\log m+m\sqrt{\log m}+k\log\log m+\textsf{Output})$ and $\mathcal{O}(n+m\sqrt{\log m})$, respectively. Using Theorem 3 instead, the total time and space will be $\mathcal{O}(n+k+\textsf{Output})$ and $\mathcal{O}(n)$, respectively. Again, our improvement is substantial (both in time and in space).
Although ephemeral edits have already been used for many other string problems, such as longest common substrings [ACI+17, ACP+20], palindrome substrings [FNI+18, FNI+19, FM21, FNI+21], Lyndon substrings [UNI+18], and substring covers [MMS+24], such a model has (to the best of our knowledge) not been defined for text indexing or pattern matching. We were thus also motivated to close this (surprising) gap in theory.
@@
In Section 2 we provide the necessary definitions and notation, as well as a few important tools that we use throughout. In Section 3, we present our solution to text indexing with ephemeral substring edits (Theorem 1). In Section 4, we present our solution to pattern matching with ephemeral block deletions (Theorem 2). In Section 5, we present our solution to pattern matching with ephemeral substring edits (Theorem 3). We conclude this paper in Section 6 with some future proposals.